Applying The Principles

you are selling your 1996 Mustang. You hav already spent $1000 on repairs. At the last time minute , the transmission dies. you can pay $600 to have it required, or sell the car "as it is. "

In each of the following scenarios, should you have the transmission required? Explain

A) Blue book value is $6500 if transmission works, $5700 if it doesen't

B) Blue book is $6000 if transmission works, $5500 if it doesnt't


1
Expert's answer
2021-02-24T15:04:17-0500

A) If transmission is repaired

Book value=$6500

Cost incurred on repairs=1000+600= $1,600

Profit= 6500-1600= $4,900

If transmission is not repaired

Book value=$5700

Cost incurred on repairs= $1000

Profit= 5700-1000 = $4,700

So, there will be more profits if the transmission is repaired, therefore, I will repair it.

B) In this case, if transmission is repaired

Profit=6000-1600= $4,600

If transmission is not repaired

Profit= 5500-1000= $4,500